Please help me with this question :
A and B are independent events. Find the missing probability :
P(A)= 3/4, P(A and B) = 1/2. find P(B)

Answer by dfrazzetto(283)


For independent events:
P(A and B) = P(A)*P(B) so:
1/2 = (3/4)P(B)



P(B) = (1/2) / (3/4) = 4/6 = 2/3